Main INDEX
Monthly INDEX
PREV
NEXT
User name jpchen
Log entry time 02:07:47 on February 7,2003
Entry number 95132
keyword=target work and status
We (Vince, Patricia and I) replaced the broken Duke with the new Exodus.
Also replaced the optics target ladder with all seven 12C foils (old 7 foils were all gone). Vince did Q-curve measurement before the cell replacement, and also
replaced the oven window for transver laser entrance.
There were a few problems:
1) The optics target ladder could not pass through the pick-up coil. While moving,
it hit the screw on the pick-up coil, cause the 12C foils to hit the glass cells. 4 out of
7 foils broke. We re-used the old optics ladder and used the all 7 foils with 4 shortened ones. The optics ladder was still too close to the left pick-up coil. We moved left pick-up coil out by ~ 1mm. Since the ladder is bulged, so we had to put the short foils in the middle to avoid the middle foils to hit the cell.
2) When moving the target up and down, there was click noise: moving down, it happened when he3 cell moved in to pickup; moving up, it happened when optics target moved into pick-up. Probably is due to the gear.
3) The reference cell is slightly off-alignment: horizontal clockwise and vertical
downwards.
We are heating up oven now. After that will do laser/optics alignment and then check
the spectro-analyzer/EPR alignment (~another 1-2 hour(s)). Then will be ready for beam.